RC Meeting, Wednesday, November 10, 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920 There is a 10:00 RC meeting today, 11/10/2021!
- Hall C to 5-pass, sharing slit with D
Hall C Status
- Controlled access for the day:
- aerogel change (Vladimir)
- general work (Heidi)
- change to 5-pass starting at 8:30 (accelerator)
- SHMS hodoscopes are off, aerogel off, HMS rotated away from beamline and SHMS ramped to 0
- Cannot operate spectrometer magnets or rotation from CH!!
- Rotate and ramp magnets for 5 pass beam run plan during day
Current Run Plan
5-pass run plan starting late day or swing: https://redmine.jlab.org/attachments/download/1317/runplan_9p9gev.pdf
- Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
- HMS Q3 insulating vacuum spikes generating annoyance alarms that take minutes to recover. https://logbooks.jlab.org/entry/3930175 Noise levels on transducers look OK, but Junaid's plot indicates Q3 is burping: https://logbooks.jlab.org/entry/3932992 .
- Serious HB LHe low liquid level alarms. https://logbooks.jlab.org/entry/3933051 Always seems to start with a subtle, simultaneous increase in level in Q1/2/3. https://logbooks.jlab.org/entry/3929152 .
- Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
- Still running with no UPS on the SHMS cryo controls - Heidi replaced, but need to check charge https://logbooks.jlab.org/entry/3937512
- HCCAM02 not functioning (lift required to replace-Brad is tracking this replacement).
- Will need a beam energy measurement
- Probably going to smallest opening angle after two shifts of beam (maybe Thursday end of day/start of swing)
- BCM calibration (Dave Mack)
- How the harp scans and carbon hole target should look (see links):
- carbon hole target beam centering procedure:
- Resetting the "ioc heartbeat epics communications" (start by clicking on the heart at upper left of the magnet controls screen
- Fall 2021 Shift Schedule
- Hall-C Live Status
- Hall-C Logbook
- MCC White Board